# A local eigenvector centrality

### v2

@@ Abstract
Eigenvector centrality is an established measure of global connectivity, from which the importance and influence of nodes can be inferred. We introduce a local eigenvector centrality that incorporates both local and global connectivity. This new measure references prominent eigengaps and combines their associated eigenspectrum, via the Euclidean norm, to detect centrality that reflects the influence of prominent community structures. In contact networks, with clearly defined community structures, local eigenvector centrality is shown to identify similar but distinct distributions to eigenvector centrality applied on each community in isolation and PageRank. Discrepancies between the two eigenvector measures highlight nodes and communities that do not conform to their defined local structures, e.g. nodes with more connections outside of their defined community than within it. While reference to PageRank’s centrality assessment enables a mitigation strategy for localisation effects inherent in eigenvector-based measures. In networks without clearly defined communities, such as city road networks, local eigenvector centrality is shown to identify both locally prominent and globally connected hubs.
@@ 1 Introduction
Eigenvector centrality is a well-established network measure that assesses node centrality in terms of global reach Bonacich1972; Bonacich2007. It is defined by the first, principal, eigenvector of a network’s adjacency matrix, $A$, that is associated with the largest eigenvalue and representative of the dominant direction of change when a vector is transformed by $A$. This measurement of a node’s ability to send and receive information across the whole network is primarily of interest when influence is dominated by a single central hub or when high global connectivity reduces the relevance of local structures. However, it is often the case for real-world networks – ranging from social and biological to technological systems – that modular, local, structures can drive global activities and function, e.g. disease spread over contact networks Clark2021; Sah2017; SalathM2010, coordination in starling flocks Clark2019, power distribution across Europe SchaubMichaelT2012, and time consensus over satellite constellations Bouis2025.
We introduce a local eigenvector centrality that considers the most topologically relevant scale for detecting node centrality, in reference to prominent eigengaps, such that the set of eigenvectors used to produce the new centrality reflects a local, community-considerate, assessment. The role of eigenvalues and their eigengaps in identifying prominent community structures can be understood through the following analogy to a physical system. Consider a spring system with two densely connected hubs and only a few connections between these two communities. By pulling such a structure in two separate directions, according to the second eigenvector, the hubs can be kept physically separate from each other at a relatively large distance. The second eigenvalue reflects this physical separation and so has a relatively large value. To pull the same structure in three distinct directions would be more difficult, with no clear division into three groups resulting in at least one of the densely connected communities being split into two sets of nodes. Therefore, the third eigenvalue would be relatively small as at least two of the three defined communities are still in close proximity. This produces a large difference between the magnitude of the second and third eigenvalues. Monitoring these differences, referred to as eigengaps or spectral gaps, provides insight into the community structure of any given graph.
@@
Eigenvector centrality was notably adapted for use in Google’s PageRank algorithm Page1999. PageRank can be conceptualised as an assessment of random walkers traversing a graph to determine the popularity of nodes as a destinations, where at each node the probability of traversing an outgoing edge is based on relative edge weightings. Eigenvector centrality has a similar conceptualisation, but instead of probabilistic flow it is the absolute edge weightings that govern the flow and assessment of node centrality. Issues arise with eigenvector centrality, referred to as localisation effects, where the centrality of densely connected hubs and isolated communities are amplified, i.e. similar to random walkers becoming trapped in an area of the graph and repeatedly visiting the same set of nodes Pastor-Satorras2016. PageRank – with its random teleportation function – and the non-backtracking (Hashimoto) matrix Arrigo2020; Pastor-Satorras2016 – that prevents return along traversed edges – have been developed in part to mitigate localisation in the calculation of centrality. Local eigenvector centrality is not developed to avoid the issue of localisation but mitigation strategies are presented herein.
@@ 2.4 Road networks
This study includes the analysis of road networks that are constructed from open-source geospatial data using OpenStreetMap (OSM). A motor-vehicle-accessible graph was extracted from OSM via the OSMnx graph_from_place function by restricting to the drive network type. The resulting graph is composed of road intersections as nodes and road links as edges. Edges are weighted according to the speed limits for given sections of road, using OSMnx’s add_edge_speeds function. Subsequently, travel times were computed, using OSMnx’s add_edge_travel_times function, yielding edge weights proportional to segment length and speed. Intersections in close spatial proximity were consolidated with OSMnx’s consolidate_intersections function with a tolerance of 0.1 and the removal of dead ends, in order to simplify the topology and remove minor junctions and roads.
The OSM-derived graph is then converted into an undirected adjacency matrix, given the largely bidirectional nature of road networks and to avoid cul-de-sacs skewing centrality results. The weighting of the adjacency are based on the inverse of the travel times, such that large weights correspond to shorter travel times. To mitigate distortions caused by extremely short links, weights are capped at a maximum value of 0.1. In the road network visualisation, speed limits are shown via a grayscale colour mapping with motorways having the highest allowable speeds and so marked in black .
@@ 3.1 Defined community structure
Figure 4: Comparison of each class eigenvector centrality with local eigenvector centrality, i=10, for the Primary School network where circle sizes are proportional to the corresponding centrality values, (b) class eigenvector centrality and PageRank difference with respect to sorted local eigenvector centrality values, and (c) boxplot of the class eigenvector centrality and PageRank (PR) differences.
@@ 3.2 Localisation mitigation
Figure 6: Comparison of (a) PageRank with local eigenvector centrality (i=5) for the High School network where circle sizes are proportional to the corresponding centrality values, (b) Local eigenvector centrality, without and with p=0.2 non-linear rescaling, difference with respect to PageRank, and (c) boxplot of the Local eigenvector centrality differences.
However, by performing a non-linear rescaling of local eigenvector centrality, $x$, as $\mathcal{N}_{p}\left(x\right)=\frac{x^{\circ p}}{\sum{x^{\circ p}}}$, where $x^{\circ p}$ is an element-wise Hadamard power, a similar distribution can be obtained for the local eigenvector centrality ($p=0.2$). The minimum power was determined by minimising the Euclidean distance, $d$, with respect to PageRank, $z$. For $p=0.2$, the distance $d(\mathcal{N}_{0.2},z)=17.1$ is similar to the Primary School comparison and much smaller than with the initial centrality $d(\mathcal{N}_{1},z)=264.4$. The main area of discrepancy after non-linear rescaling remains the high centrality values, as in Figure 4b. There is a clear rationale for discrepancies between class and local eigenvector centrality. The variation between PageRank and local eigenvector centrality is affected by differences in assessment between probabilistic and absolute flow, as well as the teleportation function. The latter promoting a more even distribution of centrality, but the largest differences are still found in high centrality values when teleportation is not enabled when applying PageRank.
@@ 3.3 Undefined community structure
High centrality values are denoted for the city of Glasgow, UK, in Figure 7 where the largest eigengap $k=5$ indicates the lack of global connectivity. The local eigenvector centrality assessment highlights five areas of high centrality that correspond to the central city centre, two prominent junctions in the city’s west end (Anniesland, Maryhill) and two in the southside of the city (Darnley, Dumbreck). The asymmetrical motorway layout, with greater access in the East end, likely influences the prominence of high centrality junctions on the less accessible west side of the city.
Figure 7: Local eigenvector centrality for the Glasgow road network. (a) Map of Glasgow with road speed limit denoted in grayscale (20–70 mph). Local eigenvector centrality values are displayed, which includes the nodes highlighted by eigenvector centrality. (b) Eigengap analysis showing the largest eigengap $k=5$.
From the detection of five key areas, we can infer that these are prominently connected to five prominent junctions/areas in Glasgow. Local effects from the scale of the graph, $n=12,979$, are responsible for the concentration of centrality at five sites. This effect masks the regions associated with each area/junction, where non-linear rescaling can be used to visualise the tributary junctions. The nodes identified by eigenvector centrality, indicated in Figure 7 are located at a single junction south of the city and visualised because they also form part of the local eigenvector assessment. The location of these globally central nodes further suggest a lack of global connectivity in the Glasgow road network, where this junction is away from the city centre and the convergence of the major motorway roads.
Glasgow displays a lack of centralisation due to a range of factors, including Glasgow’s growth integrating multiple villages and towns, a halted motorway redevelopment, and a lack of coordination between national and regional governance Cox2024. Other cities, such as Chicago in the USA, display far greater centralisation. In Figure 8, a single prominent community is detected for Chicago with the largest eigengap at $k=1$, i.e. Bonacich’s eigenvector centrality. Unlike the Glasgow example, an assessment of eigenvector centrality provides appropriate insights due to the convergence of highways and rigid grid layout creating a single central hub in the Old Town area of Chicago. The most prominent subsequent eigengap is found at $i=10$ with the local eigenvector centrality also visualised in Figure 8. It should be noted that the Chicago city boundaries extend beyond the visualised area; notably, the junction at the intersection of W 87th Street and S Vincennes Avenue is also highlighted by the local eigenvector centrality assessment.
Figure 8: Local eigenvector centrality for the Chicago road network. (a) Map of Glasgow with road speed limit denoted in grayscale (20–55 mph). Local eigenvector centrality values for $i=10$ are displayed alongside eigenvector centrality. (b) Eigengap analysis showing the largest eigengap $i=1$ and a prominent eigengap at $i=10$.
@@ 3.4 Differing insights from the Laplacian
In Shen2010, the normalised Laplacian was identified as the most effective for determining community presence from its eigengaps. Similarly Kojaku2024 found that the most effective machine learned embeddings for community detection closely resembled the spectral embeddings generated from the normalised Laplacian matrix. Hence, a case could be made for using the normalised Laplacian as the basis for a local eigenvector centrality. However, as discussed with respect to directed paths, Section 2.2, the Laplacian represents consensus dynamics and its eigenvectors highlight the nodes that can most effectively drive that collective response. Therefore, adapting the local eigenvector centrality assessment to employ the normalised Laplacian does not identify central locations and prominent junctions as is the case when using the adjacency in Figure 7 and 8. Instead, for the Glasgow road network, the Laplacian-based local eigenvector centrality shown in Figure 9 attributes the highest centrality to the most isolated communities in the graph. These communities are located around the edge of the network, and often only accessible through a single road bottleneck. With the Laplacian representing consensus dynamics, this high centrality is an acknowledgment that isolated communities require local leadership due to their lack of global connectivity Clark2019. This does not translate to the notion of centrality as detecting highly connected regions of the graph that receive significant traffic flow.
Figure 9: Local eigenvector centrality based on the normalised Laplacian of the Glasgow road network. (a) Map of Glasgow with road speed limit denoted in grayscale (20–70 mph) and Local eigenvector centrality values displayed. (b) Eigengap analysis showing the largest eigengap $k=31$.
@@ 4 Discussion
Centrality is always considered within a bounded sphere of influence, usually clearly defined by the extent of the network. In complex networks the topology is often not well characterised or understood; hence, the scale at which centrality is most relevant is less clear. For example, a study of the primary school’s contact network Stehl2011 noted high connectivity within the year groups that coincided with an expected increase in local connections within the classes. Without knowledge of the class structure, eigenvector centrality points towards the dominant presence of class 1B in Figure 3b with no central actors within years 4 and 5. In reality, each class and year group is likely to have central nodes within those communities. The local eigenvector centrality manages to capture centrality at both a year and class scale for this example, Figure 3c and d respectively. The centrality distributions at class-level validated through comparison with eigenvector centrality applied for each class subnetwork in isolation, see Figure 4.
@@
An eigengap assessment is a key step in determining a local eigenvector centrality. As noted by Shen2010; VonLuxburg2007, the presence of the largest eigengap at an index greater than 1 indicates prominent community structure. The Primary School example in Figure 3 demonstrates how a more nuanced picture of community structure emerges by investigating multiple prominent eigengaps that respond to known community structures. The largest eigengap is located at $k=1$, but prominent eigengaps are also found at $i=5$ and $i=10$ that reveal centrality distributions that respond to the year groups and classes.
@@
The Laplacian matrix, and particularly the normalised Laplacian, is often promoted as the most reliable representation for eigengap-based community detection and clustering Shen2010; VonLuxburg2007. However, a Laplacian-based centrality is influenced to a greater extent by isolated regions in a graph, which makes it less suitable as a centrality measure. The Laplacian matrix captures consensus dynamics with the consequence that Laplacian eigenvectors identify nodes required to drive a collective response Clark2019; Punzo2016. The nodes and communities with high Laplacian eigenvector entries are either highlighted due to their centrality or their isolation. The latter does not conform to the notion of centrality as a measure of highly-connected nodes that are central to prominent cycles within the network, i.e. high volume of traffic/flow. The adjacency matrix presents a better alignment to this notion of centrality by modeling the propagation of influence across the network. The adjacency is therefore better suited to calculation of local eigenvector centrality.

### v3

@@ Abstract
Eigenvector centrality is an established measure of global connectivity, from which the importance and influence of nodes can be inferred. We introduce a local eigenvector centrality that incorporates both local and global connectivity. This new measure references prominent eigengaps and combines their associated eigenspectrum, via the Euclidean norm, to detect centrality that reflects the influence of prominent community structures. In contact networks, with clearly defined community structures, local eigenvector centrality is shown to identify similar but distinct distributions to eigenvector centrality applied on each community in isolation. Discrepancies between the two eigenvector measures highlight nodes and communities that do not conform to their defined local structures, e.g. nodes with more connections outside of their defined community than within it. While reference to PageRank’s centrality assessment enables a mitigation strategy for localisation effects inherent in eigenvector-based measures. In networks without clearly defined communities, such as city road networks, local eigenvector centrality is shown to identify both locally prominent and globally connected hubs.
@@ 1 Introduction
Eigenvector centrality is a long-established network measure that assesses node centrality in terms of global reach Bonacich1972; Bonacich2007. It is defined by the first, principal, eigenvector of a network’s adjacency matrix, $A$, that is associated with the largest eigenvalue and representative of the dominant direction of change when a vector is transformed by $A$. This measurement of a node’s ability to send and receive information across the whole network is primarily of interest when influence is dominated by a single central hub or when high global connectivity reduces the relevance of local structures. However, it is often the case for real-world networks – ranging from social and biological to technological systems – that modular, local, structures can drive global activities and function, e.g. disease spread over contact networks Clark2021; Sah2017; SalathM2010, coordination in starling flocks Clark2019, power distribution across Europe SchaubMichaelT2012, and time consensus over satellite constellations Bouis2025.
We introduce a local eigenvector centrality that considers the most topologically relevant scale for detecting node centrality, in reference to prominent eigengaps, such that the set of eigenvectors used to produce the new centrality reflects a local, community-considerate, assessment. The role of eigenvalues and their eigengaps in identifying prominent community structures can be understood through the following analogy to a physical system. Consider a spring system with two densely connected hubs and only a few connections between these two communities. By pulling such a structure apart in two separate directions, according to the second eigenvector, the hubs can be kept physically separate from each other at a relatively large distance. The second eigenvalue reflects this physical separation and so has a relatively large value. To pull the same structure in three distinct directions would be more difficult, with no clear division into three groups resulting in at least one of the densely connected communities being split into two sets of nodes. Therefore, the third eigenvalue would be relatively small as at least two of the three defined communities are still in close proximity. This produces a large difference between the magnitude of the second and third eigenvalues. Monitoring these differences, referred to as eigengaps or spectral gaps, provides insight into the community structure of any given graph.
@@
Eigenvector centrality was notably adapted for use in Google’s PageRank algorithm Page1999. PageRank can be conceptualised as an assessment of random walkers traversing a graph to determine the popularity of nodes as a destination, where at each node the probability of traversing an outgoing edge is based on relative edge weightings. Eigenvector centrality has a similar conceptualisation, but instead of probabilistic flow it is the absolute edge weightings that govern the flow and assessment of node centrality. Issues arise with eigenvector centrality, referred to as localisation effects, where the centrality of densely connected hubs and isolated communities are amplified, i.e. similar to random walkers becoming trapped in an area of the graph and repeatedly visiting the same set of nodes Pastor-Satorras2016. PageRank – with its random teleportation function – and the non-backtracking (Hashimoto) matrix Arrigo2020; Pastor-Satorras2016 – that prevents return along traversed edges – have been developed in part to mitigate localisation in the calculation of centrality. Local eigenvector centrality is not developed to avoid the issue of localisation but mitigation strategies are presented herein.
@@ 2.4 Road networks
This study includes the analysis of road networks that are constructed from open-source geospatial data using OpenStreetMap (OSM). A motor-vehicle-accessible graph was extracted from OSM via the OSMnx graph_from_place function by restricting to the drive network type. The resulting graph is composed of road intersections as nodes and road links as edges. Edges have length and speed limit attributes that can be used to compute approximate travel times. If these travel times exceed what is possible, for an average car accelerating at 2 m/s2 from stationary, then the travel time, $t$, is estimated according to $t=\sqrt{\frac{2d}{a}}$ where $d$ is the distance and $a=2$ m/s2.
The OSM-derived graph is then converted into an undirected adjacency matrix, $A$, given the largely bidirectional nature of road networks and to avoid cul-de-sacs skewing centrality results. The weighting of the adjacency are based on the inverse of the travel times between nodes, $t_{ij}$ for nodes $i$ and $j$, such that large weights correspond to shorter travel times. To mitigate distortions caused by short links, such as at roundabouts and motorway on/off ramps, the weights are defined as,
$a_{ij}=\frac{1}{1+t_{ij}}.$ (12)
If the relationship was simply ${1}/{t_{ij}}$, then the function would become unbounded as $t_{ij}\rightarrow 0$.
@@ 3.1 Defined community structure
Figure 4: Comparison of each class eigenvector centrality with local eigenvector centrality, i=10, for the Primary School network where circle sizes are proportional to the corresponding centrality values, (b) class eigenvector centrality and PageRank difference with respect to sorted local eigenvector centrality values, and (c) boxplot of the class eigenvector centrality and PageRank differences.
@@ 3.2 Localisation mitigation
Figure 6: Comparison of (a) PageRank with local eigenvector centrality (i=5) for the High School network where circle sizes are proportional to the corresponding centrality values. Local eigenvector centrality, without and with p=0.2 non-linear rescaling, comparison with PageRank (PR) displayed as (b) difference area plot and (c) boxplot.
However, by performing a non-linear rescaling of local eigenvector centrality, $x$, as
$\mathcal{N}_{p}\left(x\right)=\frac{x^{\circ p}}{\sum{x^{\circ p}}}$ (13)
where $x^{\circ p}$ is an element-wise Hadamard power, a similar distribution can be obtained for the local eigenvector centrality ($p=0.2$). The minimum power was determined by minimising the Euclidean distance, $d$, with respect to PageRank, $z$. For $p=0.2$, the distance $d(\mathcal{N}_{0.2},z)=17.1$ is similar to the Primary School comparison and much smaller than with the initial centrality $d(\mathcal{N}_{1},z)=264.4$. The main area of discrepancy after non-linear rescaling remains the high centrality values, as in Figure 4b. There is a clear rationale for discrepancies between class and local eigenvector centrality. The variation between PageRank and local eigenvector centrality is affected by differences in assessment between probabilistic and absolute flow, as well as the teleportation function. The latter promoting a more even distribution of centrality, but the largest differences are still found in high centrality values when teleportation is not enabled when applying PageRank.
@@ 3.3 Undefined community structure
High centrality values are denoted for the city of Glasgow, UK, in Figure 7 where the largest eigengap $k=7$ indicates the lack of global connectivity. Localisation effects due the scale of the graph, $n=12,979$, are mitigated in Figure 7 by applying non-linear scaling to local eigenvector centrality, $p=0.2$. This expands the visible centrality distribution to more clearly identify the contributing junctions and region associated with each highly central location. The local eigenvector centrality assessment highlights seven junctions of high centrality that correspond to three junctions around the city centre (the top left city centre location is identified by Bonacich’s eigenvector centrality), one prominent junctions in the city’s west end (Maryhill), two in the southside of the city (Darnley, Battlefield) and one in the South-West (Govan) that connects the motorway to the Clyde Tunnel. The asymmetrical motorway layout, with greater access in the East end, likely influences the prominence of high centrality junctions on the less accessible west and south side of the city.
Figure 7: Local eigenvector centrality for the Glasgow road network. (a) Map of Glasgow with road speed limit denoted in grayscale (20–70 mph). Local eigenvector centrality values are displayed, which includes the nodes highlighted by eigenvector centrality. (b) Eigengap analysis showing the largest eigengap $k=7$.
The nodes identified by eigenvector centrality are adjacent to the city centre and in closest proximity to the motorway from those highlighted city centre junctions in Figure 7. Note that weighting network edges according to an inverse travel time relationship often results in motorway sections with lower weightings than close proximity junctions.
Glasgow prominent community structure is likely due to a range of factors, including Glasgow’s growth integrating multiple villages and towns, a halted motorway redevelopment, and a lack of coordination between national and regional governance Cox2024.
@@ 3.4 Differing insights from the Laplacian
In Shen2010, the normalised Laplacian was identified as the most effective for determining community presence from its eigengaps. Similarly Kojaku2024 found that the most effective machine learned embeddings for community detection closely resembled the spectral embeddings generated from the normalised Laplacian matrix. Hence, a case could be made for using the normalised Laplacian as the basis for a local eigenvector centrality. However, as discussed with respect to directed paths, Section 2.2, the Laplacian represents consensus dynamics and its eigenvectors highlight the nodes that can most effectively drive that collective response. Therefore, adapting the local eigenvector centrality assessment to employ the normalised Laplacian does not identify central locations and prominent junctions as is the case when using the adjacency in Figure 7. Instead, for the Glasgow road network, the Laplacian-based local eigenvector centrality shown in Figure 8 attributes the highest centrality to the most isolated communities in the graph. These communities are located around the edge of the network, and often only accessible through a single road bottleneck. With the Laplacian representing consensus dynamics, this high centrality is an acknowledgment that isolated communities require local leadership due to their lack of global connectivity Clark2019. This does not translate to the notion of centrality as detecting highly connected regions of the graph that receive significant traffic flow.
Figure 8: Local eigenvector centrality based on the normalised Laplacian of the Glasgow road network. (a) Map of Glasgow with road speed limit denoted in grayscale (20–70 mph) and Local eigenvector centrality values displayed. (b) Eigengap analysis showing the largest eigengap $k=37$.
@@ 4 Discussion
Centrality is always considered within a bounded sphere of influence, usually clearly defined by the extent of the network. In complex networks, the topology is often not well characterised or understood; hence, the scale at which centrality is most relevant is less clear. For example, a study of the primary school’s contact network noted high connectivity within the year groups that coincided with an expected increase in local connections within the classes Stehl2011. Without knowledge of the class structure, eigenvector centrality points towards the dominant presence of class 1B in Figure 3b with no central actors within years 4 and 5. In reality, each class and year group is likely to have central nodes within those communities. The local eigenvector centrality manages to capture centrality at both a year and class scale for this example, Figure 3c and d respectively. The centrality distributions at class-level validated through comparison with eigenvector centrality applied for each class subnetwork in isolation, see Figure 4.
@@
An eigengap assessment is a key step in determining a local eigenvector centrality. As noted by Shen2010; VonLuxburg2007, the presence of the largest eigengap at an index greater than 1 indicates prominent community structure. The Primary School example in Figure 3 demonstrates how a more nuanced picture of community structure emerges by investigating multiple prominent eigengaps that respond to known community structures. The largest eigengap is located at $k=1$, but prominent eigengaps are also found at $i=5$ and $i=10$ that reveal centrality distributions that respond to year groups and classes. Similarly, for road networks or other examples where an edge weighting must be defined in order to evaluate centrality, the eigengap distribution can vary depending on that edge weighting selection. Therefore, the largest eigengap does not provide a definitive answer on the presence of community structure. The edge weightings defined in Eq. 12 capture an inverse relationship to travel time that aligns with the notion of centrality. This relationship prevents the function from becoming unbounded when travel times between nodes are very small and produces a largest eigengap at $k=7$ for the Glasgow road network in Figure 7. By selecting a simpler inverse relationship $1/t_{ij}$ that allows for relatively large weightings on short connections, the largest eigengap appears at $k=1$ with the next most prominent eigengap clearly defined at $i=7$. Despite the differing locations of the largest eigengap, both eigengap distributions are suggestive of community structure in the Glasgow road network when the full eigengap distribution is considered.
@@
The Laplacian matrix, and particularly the normalised Laplacian, is often promoted as the most reliable representation for eigengap-based community detection and clustering Shen2010; VonLuxburg2007. However, a Laplacian-based centrality is influenced to a greater extent by isolated regions in a graph, which makes it less suitable as a centrality measure. The Laplacian matrix captures consensus dynamics with the consequence that Laplacian eigenvectors identify nodes required to drive a collective response Clark2019; Punzo2016. The nodes and communities with high Laplacian eigenvector entries are either highlighted due to their centrality or their isolation. The latter does not conform to the notion of centrality as a measure of highly-connected nodes that are central to prominent cycles within the network, i.e. high volume of traffic/flow. The adjacency matrix presents a better alignment to this notion of centrality by modelling the propagation of influence across the network. The adjacency is therefore better suited to calculation of local eigenvector centrality.
@@ 5 Conclusion
6 Data availability
Code and example datasets used in this study are currently available in a GitHub repository: https://github.com/RuaridhClark/Local_eigenvector_centrality. The repository contains MATLAB and Colab notebooks, as well as the datasets analysed in the paper. The code and data will be archived on Zenodo following paper acceptance to obtain a permanent DOI.